15.Determine whether f: R to R, defined as f (x) = −3x + 4 is a bijection. Is f invertible, and if it is, 

what is its inverse?

16.Find the inverse of f (x) =

𝑥+1

𝑥+2

, on a suitable subset of R.


1
Expert's answer
2021-12-22T09:38:46-0500

15.

Let "x_1, x_2 \in \R" and "f(x) = -3x+4."

If "f(x_1)=f(x_2)=>-3x_1+4=-3x_2+4"

"=>-3x_1=-3x_2=>x_1=x_2."

If "x_1\not=x_2=>f(x_1)\not=f(x_2)."

Then

"f(x_1)=f(x_2)<=>x_1=x_2, x_1, x_2\in \R"

The function "f(x)=-3x+4" is one-to-one.


"\forall y\in\R, y=-3x+4, \exist x=-\dfrac{1}{3}y+\dfrac{4}{3}, x\in \R"

The function "f(x)=-3x+4" is onto.


The function "f(x)=-3x+4" is one-to-one and is onto.

Therefore the function "f(x)=-3x+4" is a bijection. Therefore the function "f(x)=-3x+4" is invertibele


"f(x)=-3x+4"

Replace "f(x)" with "y"


"y=-3x+4"

Switch "x" and "y"


"x=-3y+4"

Solve for "y"


"y=-\dfrac{1}{3}x+\dfrac{4}{3}"

Replace "y" with "f^{-1}(x)"


"f^{-1}(x)=-\dfrac{1}{3}x+\dfrac{4}{3}"

16.


"f(x)=\dfrac{x+1}{x+2}"

Domain: "(-\infin, -2)\cup (-2, \infin)"

Replace "f(x)" with "y"


"y=\dfrac{x+1}{x+2}"

Switch "x" and "y"


"x=\dfrac{y+1}{y+2}"

Solve for "y"


"xy+2x=y+1"


"y=-\dfrac{2x-1}{x-1}"

Replace "y" with "f^{-1}(x)"


"f^{-1}(x)=-\dfrac{2x-1}{x-1}"